find the solution

x − 3y = −5
4x + 3y = 25

Answer by ikleyn(52788)

.
find the solution

x - 3y = -5
4x + 3y = 25
--------------------------------------

 x  - 3y  = -5    (1)
4x  + 3y  = 25    (2)

Add the two equations, (1) and (2) (both sides).
Then the terms with 3y will cancel each other, and you will get a single equation for x:

5x = -5 + 25,   or

5x = 20.

Do you know what to do next?

Please complete the solution yourself.